Potential Field Part A The lines show the equipotential contours in the plane of three point charges, Q1, Q2, and Q3. The values of the potentials are in kV as indicated for the +5, 0, and -5 kV contours. The positions of the charges are indicated by the dots 121 olta cm The letters are on the equipotential contours True: The electric field at i is stronger than at j False: The force on a proton at g points to the bottom of the page True: Q1 is a negative charge True: Charge Q3 has the largest magnitude of all False: The electric field at k is zero False: Charge Q2 is the largest negative charge You are correct. Your receipt no. is 152-867 Previous Tries Part B Calculate the work required to move a charge of -0.63x10C from i to b 3.78x10-9 j You are correct. Your receipt no. is 152-5806 Previous Tries Part C Calculate the size of the electric field at g. 2.24x105 vV/m You are correct. Your receipt no. is 152-8641 Previo Part D Calculate the size of the force on a charge of 6.40x1019 cat k 1.06×10-13 N You are correct. Your receipt no. is 152-4766 Previous Tries Part E Calculate the size of Q3. The magnitudes of the three charges are in the exact ratios of 1 to 2 to 3 ubmit Ansmor Tries 0/6
Answer #1

Solution)

Part E) Here, V1 + V2 + V3= 5000

Given, Q1:Q2:Q3= 1:2:3 (for point charges 2, 1 and 3 respectively)

So, Q1=-2Q , Q2= -Q , Q3= 3Q

-2kQ/r1 -kQ/r2 +3kQ/r3 =5000

So, Q=5000/k(-2/r1-1/r2+3/r3)

Now, we know, Q3=3Q (so, multiply value of Q by 3)

======

Good luck!:)